Form each of the following:
1. a linear equation in one variable
2. a linear equation in two variaables
3. a quadratic equation
4. a polynomial of three terms
5. an exponential function
6. a logarithmic function

Answer by stanbon(75887)

Form each of the following:
1. a linear equation in one variable
y = 6
-----------------------------------------------
2. a linear equation in two variables
x + y = 6
-----------------------------------------------
3. a quadratic equation
y = x^2 + 6x + 6
-----------------------------------------------
4. a polynomial of three terms
x^2 + 6x + 6
-----------------------------------------------
5. an exponential function
y = 6^x
--------------------------------------------------
6. a logarithmic function
y = 6^x
====================================================
Cheers,
Stan H.